OpenStudy (anonymous):

Mindy wants to solve the following system using the elimination method: y = x − 4 y = 2x − 3 What number should the equation y = x − 4 be multiplied by to eliminate x? A) -1 B)_2 C) -4 D)-6

OpenStudy (anonymous):

@GIL.ojei

OpenStudy (anonymous):

its 2 not -2

OpenStudy (anonymous):

is there 2 in the option?

OpenStudy (anonymous):

the option is -2 @GIL.ojei

OpenStudy (anonymous):

its -2

OpenStudy (anonymous):

ok thanks @GIL.ojei

OpenStudy (texaschic101):

y = x - 4 -- put in Ax + By + C form -x + y = 4 x - y = 4 y = 2x - 3 -- put in Ax + By = C -2x + y = -3 2x - y = 3 so now we have : x - y = 4 2x - y = 3 to eliminate the x term, you need to multiply first equation by -2